# Inverse problems with experiment-guided AlphaFold

### v1

@@ 2 Contributions
Hence, we introduce Density-guided AlphaFold3, which guides AlphaFold3-generated structural ensembles to be faithful to experimental electron density maps. Density-guided AlphaFold3 renders structures that are consistently more faithful to the observed electron density maps than unguided AlphaFold3İn some cases, the guided structure outperforms PDB-deposited structure’s faithfulness to the density (Table A3). Additionally, guided structures capture structural heterogeneity more accurately than AlphaFold3 (Figures 2, 3). In some cases, guided structures capture the structural heterogeneity that PDB-deposited structures fail to model (Figure 5). Lastly, we are able to leverage the strong prior learned by AlphaFold3 to generate density-faithful ensembles in a fraction of the time required by conventional X-ray crystallography pipelines (Adams et al., 2010). In our opinion, this advancement not only improves the accuracy of computational structural modeling but also has the potential to automate workflows for crystallographers.
@@ 3.1 Crystallographic electron densities
We denote the observed electron density map as $F_{\mathrm{o}}:\mathbb{R}^{3}\to\mathbb{R}$ without explicitly distinguishing between the continuous map and its discretized version. Given $F_{\mathrm{o}}$ and a protein structure ensemble $\bm{\mathcal{X}}$, the log-likelihood $\log p(F_{\mathrm{o}}\mid\bm{\mathcal{X}},\mathbf{a})$ quantifies the agreement between the experimental data and the electron density predicted by the ensemble. To compute this likelihood, we calculate the predicted electron density $F_{\mathrm{c}}(\mathbf{X})$ for each structure $\mathbf{X}$ in the ensemble (see Appendix A.3 for details). The resulting log-likelihood is given by
@@ 3.2 Nuclear Overhauser effect restraints
Interatomic distances are usually measured in NOE spectroscopy (NOESY) experiments, which consist of a two-dimensional correlation spectrum. The proximity between two atoms is evidenced by a correlation peak at a position in the spectrum that correspond to the two resonance frequencies of the spins. The intensity of the NOE signals is proportional to the temporal average of the inverse sixth power of the interatomic distance and can be used to estimate spatial constraints within the protein. NMR signal intensities depend on several other factors, which obscures the dependency on the inverse sixth power of the distance to extent. While quantitative distances can be obtained (Vögeli, 2014), it is common to use the NOE signal intensity only semi-quantitatively. With some abuse of NMR physics, we henceforth assume that the distance average is observed directly, and define the NOE constraints as a set $D=\{(\underline{d}_{ij},\overline{d}_{ij}):(i,j)\in\mathcal{P}\}$ of pairs of lower and upper bounds on the ensemble average, $d_{ij}(\bm{\mathcal{X}})=\frac{1}{n}\sum_{k=1}^{n}d_{ij}(\mathbf{X}^{k})$ of the distances $d_{ij}(\mathbf{X}^{k})=\|\mathbf{x}_{i}^{k}-\mathbf{x}_{j}^{k}\|$ between pairs of atoms $i,j$ in individual structures $\mathbf{X}^{k}$. We define the log-likelihood as
@@
where $[x]_{+}=\max(x,0)$. Additional computational details and a discussion of limitations are provided in App. A.9.
@@ 3.3 Substructure conditioning
Note that unlike the previously discussed forward models, this term is separable with respect to the individual ensemble constituents.
@@ 4.1 Guiding AlphaFold3
Figure 5: Crystallographic observations of a pair of homologous proteins (PDB: 4NE4 at $1.73$Å resolution and 5TEU at $1.62$Å). The pair features a distinct turn conformation at a corresponding site in residues 73-76 in 4NE4 and 91-94 in 5TEU despite identical local amino acid sequences and contacts with the environment (A). AlphaFold3 predicts the conformation of 4NE4 in both sequences mispredicting the conformation of 5TEU (B). Electron density-guided AlphaFold3 corrects these predictions by producing ensembles fitting well into the observed electron densities $F_{\mathrm{o}}$ of both structures depicted as the $0.3$ [e-3]-isosurfaces (C). The ensemble predicted for 5TEU exhibits conformation heterogeneity with two flipped states of the carbonyl oxygen in A82 (highlighted in red) and a flexible side chain in A82 (dark green) better explaining the $F_{\mathrm{o}}$ (D).
@@ 4.3 Ensemble filtering using matching pursuit
Post relaxation, we employ a matching pursuit-based approach (Mallat & Zhang, 1993) to greedily select a subset of the relaxed ensemble, $\bm{\mathcal{X}}_{\mathcal{I}}=\{\mathbf{X}^{k}:k\in\mathcal{I}\}$, that best fits the observation $\mathbf{y}$. Starting with $\mathcal{I}=\emptyset$, during every iteration of the matching pursuit algorithm, we seek to maximize $\log p(\mathbf{y}\mid\bm{\mathcal{X}}_{\mathcal{I}\cup\{k\}},\mathbf{a})$ over all $k\notin\mathcal{I}$. The optimal element $k$ is then added to the support set $\mathcal{I}$ and the process is repeated until the likelihood no longer increases or the maximum allowed ensemble size $n_{\text{max}}$ is reached. A detailed explanation of the matching pursuit-based ensemble filtering procedure, along with its pseudocode, is provided in Appendix A.5.
@@ 5 Modeling crystallographic ensembles
Heterogeneity in equivalent environments. The second case we explore is a more local and subtle version of the former, highlighting that the same amino acid segment in distinct but homologous proteins can adopt different conformations, even when the immediately surrounding protein environment is equivalent (and, thus, the local segments have identical contact networks). A key component of AlphaFold3 prediction is based on coevolutionary signals from multiple sequence alignment (MSA) of homologous proteins in order to discern contact maps (which, in the form of MSA embedding, condition the diffusion model as depicted in Figure 1). It is, therefore, unsurprising then that homologous proteins harboring regions of identical amino acid sequence embedded into an equivalent spatially adjacent amino acid environment are predicted to have identical local conformation even from distinct protein sequences. As an illustration, we show in Figure 5 a pair of such homologous proteins (PDB 4NE4 and 5TEU) featuring a sequence of $4$ amino acids in two distinct turn conformations. While the 5TEU conformation variant is mispredicted by AlphaFold3, adding electron density guidance reproduces both variants faithfully. We also observe a more heterogeneous ensemble generated in the latter case, including flipped carbonyl oxygen and sidechain flexibility that better explain the density. A quantitative evaluation of this and additional homologous pairs is presented in Table A2.
Figure 6: Comparison of conformational flexibility in ubiquitin ensembles predicted by AlphaFold3 and our i.i.d and non-i.i.d NOE-guided method. The N-H order parameter $\mathrm{S}^{2}$ is calculated on the ensembles and compared to the experimental NMR observables. For guided predictions, plotted are the mean and standard deviation on $5$ independent runs. Green and purple shadings indicate $\alpha$-helices and $\beta$-strands as predicted by DSSP (A). AlphaFold3 exhibits the lowest correlation to the experimental measurement ($r=0.47$), while guiding AlphaFold3 with NOE measurements improves it to $r=0.52$ for i.i.d guidance and $r=0.72$ with non-i.i.d. ensembles (B). For two lysines at site 11 (entrance into the second $\beta$-strand from a flexible loop) and 29 (middle of $\alpha$-helix) highlighted in (C), guided ensembles exhibit more variability of the amide N-H bond direction in the flexible turn, while showing less variability in the inflexible helix, in agreement with the experimental observation (D).
@@ 6 Modeling NMR ensembles
NMR exploits the magnetic resonance of atomic nuclei to probe into protein structure and dynamics. NOEs acquired from solution-state NMR, in particular, provide distances between atoms averaged over the ensemble of molecules in the sample and over time scales up to milliseconds. NOE-derived distances, thus, comprise the conformational heterogeneity.
Status quo. NMR structures are determined by integrating molecular dynamics (MD) simulations with NMR-derived restraints, using biomolecular force fields (Wang et al., 2004; Schwieters et al., 2006; Güntert, 2004). However, since NMR observables inherently reflect ensemble-averaged measurements, simulating single conformers often leads to mode collapse, producing rigid ensembles that poorly capture true conformational dynamics (Figure 4 in Lindorff-Larsen et al. (2005)). To address this, ensemble-based MD approaches—pioneered by (Lindorff-Larsen et al., 2005) and (Lange et al., 2008)—simulate multiple conformers simultaneously to satisfy experimental restraints. While effective, these methods remain computationally prohibitive, requiring days even for small systems like the $76$-residue ubiquitin. While AlphaFold has revolutionized structure prediction, its training on static X-ray crystallography data biases its ensembles toward rigid conformations, failing to capture conformational heterogeneity or satisfy NMR-derived restraints (Figure 7 and Table A4). This limits its utility for modeling protein dynamics. In what follows, we demonstrate that NOE-guided AlphaFold3 recovers ensembles that (i) rigorously obey NOE distance restraints; (ii) reproduce experimentally observed flexibility; and (iii) achieve this in minutes, overcoming the computational bottleneck of traditional ensemble methods.
@@
Figure 7: Conformation ensembles generated for six proteins using AlphaFold3 (first row) and the proposed NOE guidance (second row). Ensembles are visualized in blue overlaid on corresponding NMR structures solved from the same NOESY data. PDB identifiers are indicated above each structure. The numbers below report the percentage of violated constraints and the median violation.
Peptides. We used the benchmark from McDonald et al. (2023) from which we selected $20$ peptides worst predicted by AlphaFold3, for which the NMR structures result violate less than $10\%$ distance restraints. This resulted in three structures, namely, 1DEC, 2LI3, 3BBG (Figure 7), for which AlphaFold3 produces a partially wrong fold than what is suggested by NMR measurements. This is visually depicted in Figure 7, and is evidenced by the violation of restraints of AlphaFold3 baseline presented in Table A4. We observed that NOE-guided AlphaFold3, both produces ensembles that obey $>15\%$ more restraints and an order of magnitude lower error (see Table A4) and fixes the misprediction made by unguided AlphaFold3 (Figure 7).
100 NMR spectra database. We evaluated the proposed method on $9$ proteins ($55-102$ residues) from a recently compiled database of NMR structures (Klukowski et al., 2024). While AlphaFold3 largely retains the correct fold for these systems, its baseline predictions exhibit, on average,  8% more NOE restraint violations than the deposited NMR structures (Table A4). In all cases, NOE-guided AlphaFold reduced violations, and in half of the cases, the guided ensembles even outperformed PDB-deposited NMR structures in restraint compliance. Visual alignment with NMR ensembles (Figure 7) and quantitative agreement in $\rho$-RMSF (Table A4) confirm that guidance recovers conformational heterogeneity consistent with NMR structures. Increasing the guidance scale progressively enforces restraint satisfaction, as shown in Figure 4. Critically, these improvements are achieved within minutes—reaching the accuracy of MD-derived ensembles at a fraction of the computational cost.
@@ 7 Conclusion
In this paper, we presented a general methodology for guiding AlphaFold3 using experimental observables and systematically evaluated its utility in fitting structural ensembles to crystallographic and NMR data. In crystallography, the robustness and generality of our method are demonstrated by the wide range of conditions of the evaluated targets: electron density maps covering a wide range of resolutions (from sub-Å to medium-low resolution) and qualities (covering high and low B-factors), protein segments with different lengths, variable amino acid identities, different secondary structure contexts, and AlphaFold3 mispredictions of severity ranging from the subtle conformation of the carbonyl oxygen to loop regions predicted tens of Ångstroms away from the experimentally observed location. Likewise, in NMR we show a great variability of the number and quality of NOE constraints in the evaluated structures. In the future, we intend to generalize the method to jointly sampling ensembles representing multiple molecules to model protein complexes and protein-ligand interactions. We also plan to extend the model to single- and multi-particle cryoEM where handling conformational heterogeneity constitutes a major challenge.
@@ A.1 Guidance using AlphaFold3
To ensure numerical stability during guided diffusion, we apply gradient clipping (Zhang et al., 2019) to clip the guidance score. This prevents instability due to large gradients, ensuring a smooth integration of experimental constraints into the diffusion process.
A.2 AlphaFold3 model and hardware resources
Across all experiments in this paper, we used the open-sourced Protenix (Chen et al., 2025) model, a PyTorch-based (Paszke et al., 2019) reimplementation of AlphaFold3. However, for the AlphaFold3 baseline comparisons, we report predictions generated using the official AlphaFold3 weights and source code (Abramson et al., 2024). All computations were performed on NVIDIA H100 and L40S GPUs.
A.3 Calculating $F_{\mathrm{c}}$ from a protein structure

### v2

@@ 2 Contributions
Hence, we introduce Density-guided AlphaFold3, which guides AlphaFold3-generated structural ensembles to be faithful to experimental electron density maps. Density-guided AlphaFold3 renders structures that are consistently more faithful to the observed electron density maps than unguided AlphaFold3. In some cases, the guided structure outperforms PDB-deposited structure’s faithfulness to the density (Table A3). Additionally, guided structures capture structural heterogeneity better than AlphaFold3 (Figures 2, 3). In some cases, guided structures capture the structural heterogeneity that PDB-deposited structures fail to model (Figure A1). Lastly, we are able to leverage the strong prior learned by AlphaFold3 to generate density-faithful ensembles in a fraction of the time required by conventional X-ray crystallography pipelines (Adams et al., 2010) (Table A5). In our opinion, this advancement not only improves the accuracy of computational structural modeling but also has the potential to automate workflows for crystallographers.
@@ 3.1 Crystallographic electron densities
We denote the observed electron density map as $F_{\mathrm{o}}:\mathbb{R}^{3}\to\mathbb{R}$ without explicitly distinguishing between the continuous map and its discretized version. Given $F_{\mathrm{o}}$ and a protein structure ensemble $\bm{\mathcal{X}}$, the log-likelihood $\log p(F_{\mathrm{o}}\mid\bm{\mathcal{X}},\mathbf{a})$ quantifies the agreement between the experimental data and the electron density predicted by the ensemble. To compute this likelihood, we calculate the predicted electron density $F_{\mathrm{c}}(\mathbf{X})$ for each structure $\mathbf{X}$ in the ensemble (see Appendix A.2, A.6 for details). The log-likelihood is given by
@@ 3.2 Nuclear Overhauser effect restraints
Interatomic distances are usually measured in NOE spectroscopy (NOESY) experiments, which consist of a two-dimensional correlation spectrum. The proximity between two atoms is evidenced by a correlation peak at a position in the spectrum that correspond to the two resonance frequencies of the spins. The intensity of the NOE signals is proportional to the temporal average of the inverse sixth power of the interatomic distance and can be used to estimate spatial constraints within the protein. NMR signal intensities depend on several other factors, which obscures the dependency on the inverse sixth power of the distance to extent. While quantitative distances can be obtained (Vögeli, 2014), it is common to use the NOE signal intensity only semi-quantitatively. With some abuse of NMR physics, we henceforth assume that the distance average is observed directly, and define the NOE constraints as a set $D=\{(\underline{d}_{ij},\overline{d}_{ij}):(i,j)\in\mathcal{P}\}$ of pairs of lower and upper bounds on the ensemble average, $d_{ij}(\bm{\mathcal{X}})=\frac{1}{n}\sum_{k=1}^{n}d_{ij}(\mathbf{X}^{k})$ of the distances $d_{ij}(\mathbf{X}^{k})=\|\mathbf{x}_{i}^{k}-\mathbf{x}_{j}^{k}\|$ between pairs of atoms $i,j$ in individual structures $\mathbf{X}^{k}$. The log-likelihood is given by,
@@
where $[x]_{+}=\max(x,0)$. Additional computational details, noise model, and limitations are provided in App. A.2, A.12.
@@ 3.3 Substructure conditioning
Note that unlike the previously discussed forward models, this term is separable with respect to the individual ensemble constituents. See A.2 for the noise model.
@@ 4.3 Ensemble filtering using matching pursuit
Post relaxation, we employ a matching pursuit-based approach (Mallat & Zhang, 1993) to greedily select a subset of the relaxed ensemble, $\bm{\mathcal{X}}_{\mathcal{I}}=\{\mathbf{X}^{k}:k\in\mathcal{I}\}$, that best fits the observation $\mathbf{y}$. Starting with $\mathcal{I}=\emptyset$, during every iteration of the matching pursuit algorithm, we seek to maximize $\log p(\mathbf{y}\mid\bm{\mathcal{X}}_{\mathcal{I}\cup\{k\}},\mathbf{a})$ over all $k\notin\mathcal{I}$. The optimal element $k$ is then added to the support set $\mathcal{I}$ and the process continues until the likelihood no longer increases or the maximum allowed ensemble size $n_{\text{max}}$ is reached. A detailed explanation of the matching pursuit-based ensemble filtering procedure, along with its pseudocode, is provided in Appendix A.8.
@@ 5 Modeling crystallographic ensembles
Figure 5: Comparison of conformational flexibility in ubiquitin ensembles predicted by AlphaFold3 and our i.i.d and non-i.i.d NOE-guided method. The N-H order parameter $\mathrm{S}^{2}$ is calculated on the ensembles and compared to the experimental NMR observables. For guided predictions, plotted are the mean and standard deviation on $5$ independent runs. Green and purple shadings indicate $\alpha$-helices and $\beta$-strands as predicted by DSSP (Frishman & Argos, 1995) (A). AlphaFold3 exhibits the lowest correlation to the experimental measurement ($r=0.47$), while guiding AlphaFold3 with NOE measurements improves it to $r=0.52$ for i.i.d guidance and $r=0.72$ with non-i.i.d. ensembles (B). For two lysines at site 11 (entrance into the second $\beta$-strand from a flexible loop) and 29 (middle of $\alpha$-helix) highlighted in (C), guided ensembles exhibit more variability of the amide N-H bond direction in the flexible turn, while showing less variability in the inflexible helix, in agreement with the experimental observation (D).
@@
Heterogeneity in equivalent environments. The second case we explore is a more local and subtle version of the former, highlighting that the same amino acid segment in distinct but homologous proteins can adopt different conformations, even when the immediately surrounding protein environment is equivalent (and, thus, the local segments have identical contact networks). A key component of AlphaFold3 prediction is based on coevolutionary signals from multiple sequence alignment (MSA) of homologous proteins in order to discern contact maps (which, in the form of MSA embedding, condition the diffusion model as depicted in Figure 1). It is, therefore, unsurprising then that homologous proteins harboring regions of identical amino acid sequence embedded into an equivalent spatially adjacent amino acid environment are predicted to have identical local conformation even from distinct protein sequences. As an illustration, we show in Figure A1 a pair of such homologous proteins (PDB 4NE4 and 5TEU) featuring a sequence of $4$ amino acids in two distinct turn conformations. While the 5TEU conformation variant is mispredicted by AlphaFold3, adding electron density guidance reproduces both variants faithfully. We also observe a more heterogeneous ensemble generated in the latter case, including flipped carbonyl oxygen and sidechain flexibility that better explain the density. A quantitative evaluation of this and additional homologous pairs is presented in Table A2.
@@
An extended analysis is presented in Table A9, with additional $15$ proteins where regions up to 22 residues in some cases (5V2M & 6E2S) are optimized. In all cases, cosine similarity between observed and calculated electron density maps matched or exceeded that of the PDB entries. Similar results were observed for polypeptide chains.
@@ 6 Modeling NMR ensembles
NMR exploits the magnetic resonance of atomic nuclei to probe protein structure and dynamics. NOEs acquired from solution-state NMR, in particular, provide distances between atoms averaged over the ensemble of molecules in the sample and over time scales up to milliseconds. NOE-derived distances, thus, comprise the conformational heterogeneity.
Status quo. NMR structures are determined by integrating molecular dynamics (MD) simulations with NMR-derived restraints, using biomolecular force fields (Wang et al., 2004; Schwieters et al., 2006; Güntert, 2004). However, since NMR observables inherently reflect ensemble-averaged measurements, simulating single conformers often leads to mode collapse, producing rigid ensembles that poorly capture true conformational dynamics (Figure 4 in Lindorff-Larsen et al. (2005)). To address this, ensemble-based MD approaches—pioneered by (Lindorff-Larsen et al., 2005) and (Lange et al., 2008)—simulate multiple conformers simultaneously to satisfy experimental restraints. While effective, these methods remain computationally prohibitive, requiring days even for small systems like the $76$-residue ubiquitin. While AlphaFold has revolutionized structure prediction, its training on static X-ray crystallography data biases its ensembles toward rigid conformations, failing to capture conformational heterogeneity or satisfy NMR-derived restraints (Figure A7 and Table A4). This limits its utility for modeling protein dynamics. In what follows, we demonstrate that NOE-guided AlphaFold3 recovers ensembles that (i) rigorously obey NOE distance restraints; (ii) reproduce experimentally observed flexibility; and (iii) achieve this in minutes, overcoming the computational bottleneck of traditional ensemble methods.
@@
Peptides. We used the benchmark from McDonald et al. (2023) from which we selected $20$ peptides worst predicted by AlphaFold3, for which the NMR structures result violate less than $10\%$ distance restraints. This resulted in three structures, namely, 1DEC, 2LI3, 3BBG (Figure A7), for which AlphaFold3 produces a partially wrong fold than what is suggested by NMR measurements. This is visually depicted in Figure A7, and is evidenced by the violation of restraints of AlphaFold3 baseline presented in Table A4. We observed that NOE-guided AlphaFold3, both produces ensembles that obey $>15\%$ more restraints and an order of magnitude lower error (see Table A4) and fixes the misprediction made by unguided AlphaFold3 (Figure A7).
100 NMR spectra database. We evaluated the proposed method on $9$ proteins ($55-102$ residues) from a recently compiled database of NMR structures (Klukowski et al., 2024). While AlphaFold3 largely retains the correct fold for these systems, its baseline predictions exhibit, on average,  8% more NOE restraint violations than the deposited NMR structures (Table A4). In all cases, NOE-guided AlphaFold reduced violations, and in half of the cases, the guided ensembles even outperformed PDB-deposited NMR structures in restraint compliance. Visual alignment with NMR ensembles (Figure A7) and quantitative agreement in $\rho$-RMSF (Table A4) confirm that guidance recovers conformational heterogeneity consistent with NMR structures. Increasing the guidance scale progressively enforces restraint satisfaction, as shown in Figure 4. Critically, these improvements are achieved within minutes (Table A6) – reaching the accuracy of MD-derived ensembles at a fraction of the computational cost.
A supplementary analysis is presented in Table A10, featuring $27$ additional NMR proteins from NMRDb (Banfi & Patiny, 2008). For each protein, the guided ensemble better satisfies NOE restraints than unguided AlphaFold3 structures, and sometimes, even outperforms the corresponding NMR PDB structure.
7 Related Work
This work presents the first application of guided sampling with AlphaFold3 to resolve protein structural ensembles using crystallographic and NMR experimental data. Prior approaches have attempted similar goals but exhibit key limitations. Toth-Petroczy et al. (2016) uses a Potts model (Wu, 1982) to predict structured regions in intrinsically disordered proteins from evolutionary data. However, they do not leverage experimental measurements. Fadini et al. (2025) fit a single conformer to density maps by optimizing MSA contact maps but cannot model conformational heterogeneity. Liu et al. (2024) use str2str protein diffusion model (Lu et al., 2014) to generate cryo-EM–guided structural ensembles. While related, their focus is on a different modality. Maddipatla et al. (2024) fits ensembles to crystallographic density maps using Chroma (Ingraham et al., 2023) and captures multiple conformers to some extent. However, it fails to capture conformational heterogeneity over a long residue range due to Chroma’s hierarchical formulation. Also, Chroma’s sequence conditioning is only promoted, limiting enforcement of NMR distances. Finally, Levy et al. (2024) uses Chroma as their diffusion model and inherit similar limitations. Additionally, they use synthetic data instead of raw experimental observations.
@@ 8 Conclusion
In this paper, we present a general methodology for guiding AlphaFold3 using experimental observables and systematically evaluated its utility in fitting structural ensembles to crystallographic and NMR data. In crystallography, the robustness of our method demonstrated by the wide range of conditions of the evaluated targets: electron density maps covering a wide range of resolutions (from sub-Å to medium-low resolution) and qualities (covering high and low B-factors), protein segments with different lengths, variable amino acid identities, different secondary structure contexts, and AlphaFold3 mispredictions of severity ranging from the subtle conformation of the carbonyl oxygen to loop regions predicted tens of Ångstroms away from the experimentally observed location. Likewise, in NMR we show a great variability of the number and quality of NOE constraints in the evaluated structures. In the future, we intend to generalize the method to jointly sampling ensembles representing multiple molecules to model protein complexes and protein-ligand interactions. We also plan to extend the model to single- and multi-particle cryoEM where handling conformational heterogeneity constitutes a major challenge.
Acknowledgments
This work was supported by the Israeli Science Foundation (ISF) grant number $1834/24$. We acknowledge support from the Austrian Science Fund (FWF, grant numbers I5812-B and I6223) and the financial support of the Helmsley Fellowships Program for Sustainability and Health. This research uses resources of the Institute of Science and Technology Austria’s scientific computing cluster.
@@ A.1 Guidance using AlphaFold3
To ensure numerical stability during guided diffusion, we apply gradient clipping to clip the guidance score. This prevents instability due to large gradients, ensuring a smooth integration of experimental constraints into the diffusion process.
A.2 Noise model underlying log likelihood
This section discusses the underlying noise models $\mathcal{L}$ that lead to the log-likelihood functions in Equations 1, 3.2, and 3.
  • Density Loss: A Laplace noise model is used, where the difference between $F_{\mathrm{o}}$ and $F_{\mathrm{c}}$ is drawn from a Laplace distribution centered at zero with unit scaling. This is given by,
    $\mathcal{L}=F_{\mathrm{o}}-F_{\mathrm{c}}\sim\text{Laplace}(0,1)$ (6)
    This model, along with the Gaussian model, is used to model electron density. However, a more realistic noise model involves complex physics, as noise is introduced at the level of Fourier intensities. We will address this in future works.
  • NOE Loss: Instead of a typical noise model, a piecewise function is used to model the underlying noise function.
    =$\mathcal{L}=\begin{cases}0,&\text{if }\underline{d}_{ij}\leq d_{ij}(\bm{\mathcal{X}})\leq\overline{d}_{ij}\\ (d_{ij}(\bm{\mathcal{X}})-\underline{d}_{ij})^{2},&\text{if }d_{ij}(\bm{\mathcal{X}})<\underline{d}_{ij}\\ (d_{ij}(\bm{\mathcal{X}})-\overline{d}_{ij})^{2},&\text{if }d_{ij}(\bm{\mathcal{X}})>\overline{d}_{ij}\end{cases}$ (7)
    In Equation 7, if the pairwise distances $d_{ij}(\bm{\mathcal{X}})$ is between the bounds $\left[\underline{d}_{ij},\overline{d}_{ij}\right]$, then we assume a uniform noise distribution (00 loss). Otherwise, we assume a Gaussian-like noise distribution (quadratic).
  • Substructure Loss: A Gaussian distribution with a fixed isotropic covariance is used as the noise model.
    $\mathcal{L}=\mathbf{x}_{i}^{k}\sim\mathcal{N}(\mathbf{y}_{i},\mathbf{I})$ (8)
    Here, each atom $\mathbf{x}_{i}^{k}$ is drawn from a multivariate Gaussian distribution centered at reference atom location $\mathbf{y}_{i}$ with identity covariance matrix $\mathbf{I}$ (isotropic).
A.3 Baselines
We extend the evaluation presented in Tables A3 and A4 for both X-ray crystallography and NMR experimental observations by incorporating additional baselines, specifically AlphaFlow and ESMFlow (Jing et al., 2024). For the X-ray crystallography, we include a comparison with Chroma ensembles generated using non-i.i.d. guidance (Maddipatla et al., 2024). Table A8 presents a direct comparison between non-i.i.d. guided X-ray crystallography (which consistently outperforms other baselines in Table A3) and the newly added baselines. In all cases, non-i.i.d. guided AlphaFold3 demonstrates superior performance compared to other methods.
Similarly, Table A4 has been extended to include AlphaFlow and ESMFlow in the NMR structure benchmark, as shown in Table A7. Our results indicate that the structural ensembles generated by NOE-guided AlphaFold3 adhere more closely to experimental constraints than those produced by any other baseline. ably, in half of the cases, these ensembles show better agreement with the constraints than the deposited NMR structures resolved using molecular dynamics. It is important to note that due to the absence of explicit sequence conditioning in Chroma, the resulting ensembles diverged significantly from the true structures.
A.4 Runtime Analysis
In addition to predicting experimentally faithful protein ensembles, the proposed approach renders samples in a computationally efficient manner. As shown in Table A5, an ensemble of $16$ proteins with over $300$ residues is sampled in approximately $7$ minutes using density guidance, with minimal added latency compared to unguided AlphaFold3. Similarly, in Table A6, we generate ensembles of $32$ conformations in approximately $14$ minutes while modeling distance restraints as ensemble statistics. Our approach is significantly faster than restrained molecular dynamics methods like CYANA (Güntert, 2004), which is the current state-of-the-art technique for NMR structure determination.
A.5 AlphaFold3 model and hardware resources
Across all experiments in this paper, we used the open-sourced Protenix (Chen et al., 2025) model, a PyTorch-based (Paszke et al., 2019) re-implementation of AlphaFold3. However, for the AlphaFold3 baseline comparisons, we report predictions generated using the official AlphaFold3 weights and source code (Abramson et al., 2024). All computations were performed on NVIDIA H100 and L40S GPUs.
A.6 Calculating $F_{\mathrm{c}}$ from a protein structure